OpenStudy (anonymous):

If a 20-sided fair die with sides distinctly numbered 1 through 20 is rolled, the probability that the answer is a perfect square can be expressed as ab where a and b are coprime positive integers. What is the value of a+b?

OpenStudy (anonymous):

Do you mean probability expressed as a/b? Perfect squares from 1-20 are 1, 4, 9 and 16. So P(ans is sqaure) = 4/20 =1/5 So a=1, b=5, and a+b=6